# One-Step Diffusion-Based Image Compression with Semantic Distillation

### v1

@@ Abstract
While recent diffusion-based generative image codecs have shown impressive performance, their iterative sampling process introduces unpleasing latency. In this work, we revisit the design of a diffusion-based codec and argue that multi-step sampling is not necessary for generative compression. Based on this insight, we propose OneDC, a One-step Diffusion-based generative image Codec—that integrates a latent compression module with a one-step diffusion generator. Recognizing the critical role of semantic guidance in one-step diffusion, we propose using the hyperprior as a semantic signal, overcoming the limitations of text prompts in representing complex visual content. To further enhance the semantic capability of the hyperprior, we introduce a semantic distillation mechanism that transfers knowledge from a pretrained generative tokenizer to the hyperprior codec. Additionally, we adopt a hybrid pixel- and latent-domain optimization to jointly enhance both reconstruction fidelity and perceptual realism. Extensive experiments demonstrate that OneDC achieves SOTA perceptual quality even with one-step generation, offering over 40% bitrate reduction and 20$\times$ faster decoding compared to prior multi-step diffusion-based codecs. Code will be released later.
footnotetext: Naifu Xue and Zhaoyang Jia are visiting students at Microsoft Research Asia. This paper stems from an open-source project.
@@ 1 Introduction
The rapid growth in image data has led to increased storage and transmission costs, heightening the need for efficient, high-quality image compression techniques. While VAE-based learned image compression (LIC) [31, 63] has surpassed traditional codecs (e.g., VVC [6]) in rate-distortion (RD) performance, it often produce blurry details at low bitrates [36]. To mitigate this, some approaches [36, 39] focus on optimizing visual quality by incorporating perceptual metrics and adversarial losses to enable generative image compression. Despite these advances, such generative models tend to introduce artifacts under extreme compression, degrading the realism of reconstructed images.
@@
In this paper, we introduce OneDC (One-step Diffusion-based generative Codec), a novel framework for ultra-low bitrate image coding. OneDC integrates a latent compression module with a one-step diffusion generator: the former encodes the image into compact latent, while the latter synthesizes high-frequency details conditioned on the latent. Since the pixel-domain training is insufficient for guaranteeing perceptual quality at low bitrates [18], we adopt a hybrid-domain training strategy. Specifically, we combine a pixel-domain perceptual loss to promote fidelity with a latent-domain diffusion distillation objective [58] to enhance realism. This design allows OneDC to effectively leverage the pretrained diffusion model while achieving a balanced trade-off among compression ratio, perceptual quality, and decoding efficiency, as illustrated in Fig. 1.
Moreover, we further explore the role of semantic guidance (i.e., the input of the cross-attention layers [44]) within our framework. In one-step diffusion, such guidance is essential to compensate for the absence of multi-step refinement (See ablation in Section 4.3). While existing diffusion models have utilized text prompts for conditioning, we argue that text is suboptimal for image compression due to two limitations: (1) natural language struggles to capture fine-grained or localized visual semantics, and (2) generating high-quality captions typically requires large-scale vision-language models (e.g., the large BLIP2 [27] used in [7]), introducing substantial computational overhead. This raises a central question: Can we design more effective semantic representations than text for guiding one-step diffusion-based codecs—without incurring excessive computation cost?
Recent studies have shown that the hyperprior in VAE-based codecs can capture high-level semantic information [18]. Compared to textual prompts, hyperpriors provide more precise and spatially aligned semantic cues, making them well-suited for representing localized semantics in high-resolution images (see Fig.2). This insight motivates us to further enhance the representational capability of the hyperprior. Inspired by generative tokenizers [12], where codebooks encode rich semantic content, we propose hyperprior semantic distillation—a training strategy that transfers semantic knowledge from a pretrained tokenizer to the hyperprior through a code prediction module [64]. As illustrated in Fig.2, this approach results in reconstructions that are more semantically accurate and visually coherent.
Experiments show that OneDC achieves state-of-the-art (SOTA) performance in generative image compression, delivering over 40% bitrate reduction and 20$\times$ faster decoding compared to existing multi-step diffusion codecs. Ablation studies further validate the effectiveness of our optimization strategy and the semantic distillation method for hyperprior. Our contributions are as follows:
@@ 2 Related Work
VAE-based Generative Image Compression. VAE-based codecs [31, 63, 50, 4, 38] often produce blurry reconstructions at low bitrates, limiting their perceptual quality. To overcome this, GAN-based approaches such as Agustsson et al. [2] and HiFiC [36] train models to synthesize textures using adversarial discriminators. MS-ILLM [39] improves realism through a local likelihood model, while EGIC [22] introduces semantic-aware feedback via segmentation-conditioned discriminators. Other methods enhance compression by integrating generative tokenizers: Mao et al. [35] utilize VQGAN [12] for compression, Xue et al. [55] unify token generation and entropy coding, and Jia et al. [18] propose GLC that improves token compression via transform coding. More recently, dual-branch codecs [34, 56] have been introduced to support more compact and flexible latent representations. While these methods highlight the strong feature learning capacity of VAE-based codecs, their reconstruction quality at extremely low bitrates remains limited by model capacity and training scale.
@@
Diffusion-based Generative Image Compression. Recent works [24, 7, 28, 49, 54] have explored the use of generative priors from large-scale pretrained diffusion models in image compression, demonstrating improved perceptual realism compared to conventional VAE-based approaches. Eric et al. [24] combinepretrained diffusion with ControlNet [61] to reconstruct images from captions and compressed sketches. PerCo [7] fine-tunes diffusion models using captions and quantized features, with captions generated by a large BLIP2 model [27]. DiffEIC [28] conditions diffusion models on VAE latents, showing strong performance even without text, and their subsequent work [29] improves encoder and introduces relay fine-tuning. Relic et al. [43] treat quantization error as noise and apply denoising diffusion for reconstruction. DiffC [49] demonstrates zero-shot compression with diffusion models using reverse-channel coding. While these approaches improve realism at ultra-low bitrates, they often struggle to maintain high fidelity. Moreover, multi-step sampling in DDIM [45] introduces substantial decoding overhead, limiting their practicality. By contrast, our OneDC achieves high fidelity and realism with significantly faster decoding via one-step sampling. While prior work [28] suggests high-level semantics (e.g., text) may be optional, we find such guidance remains crucial in our one-step diffusion for image decoding task, as detailed in Section 4.3.
@@
One-step Diffusion Models. To reduce the high computational cost of diffusion models, recent efforts have focused on distilling multi-step models into efficient one-step generators. Yin et al. [59] propose Distribution Matching Distillation (DMD), which minimizes an approximate KL divergence between real and generated image distributions modeled by multi-step teacher networks. Their subsequent work, DMD2 [58], further improves generation quality by introducing a diffusion-GAN framework and enhanced training strategies. Based on DMD, Song et al. [46] introduce class-specific student models to improve performance across diverse categories. SwiftBrush series [40, 9] adopt a LoRA adaptation [17] for variational score distillation and further optimize diversity by integrating a clamped CLIP loss. Independently, Kang et al. [20] employ perceptual loss supervision and a multi-scale U-Net discriminator for one-step distillation. These models have also been extended to low-level vision tasks such as super-resolution [11, 53, 51, 21] and image restoration [15], highlighting the potential of one-step diffusion beyond generation. Inspired by these advances, we explore its application in generative image compression, a setting that remains underexplored.
@@ 3 Methodology
We propose OneDC, a one-step diffusion-based generative image compression framework that achieves high-quality reconstruction with low-latency decoding under ultra-low bitrate constraints. Section 3.1 introduces the coding pipeline of our method, consisting of a latent compression module that encodes the input image into compact latent representation and a one-step diffusion generator that synthesizes reconstructions with realistic details. Given the critical role of semantic guidance in our one-step diffusion and the limitations of textual prompts in compression scenarios, Section 3.2 introduces the use of the hyperprior as an alternative conditioning signal. To further tap the potential of the hyperprior, we propose a semantic distillation strategy that transfers knowledge from a pretrained generative tokenizer to the hyperprior codec, thereby improving semantic accuracy. Finally, Section 3.3 presents our training scheme, which combines pixel-domain and latent-domain objectives to jointly optimize reconstruction fidelity and perceptual realism.
@@ 3.1 Framework Overview
Fig. 3 shows the overview of our OneDC framework. On the encoder side, an analysis transform encodes the input $x$ into a compact latent representation $y=g_{a}(x)$, which is then quantized to $\hat{y}$. A hyper encoder further processes the latent into a hyperprior $z=h_{enc}(y)$, quantized as $\hat{z}$. Then, the entropy model takes decoded hyperprior context $c_{h}=h_{ctx}(\hat{z})$ and spatial context $c_{s}$ to predict the distribution of $\hat{y}$, enabling bitrate estimation during training and entropy coding during inference.
@@
On the decoder side, the received $\hat{y}$ is passed through a synthesis transform to produce an initial latent $\tilde{y}_{t}=g_{s}(\hat{y})$ for diffusion generation. Given the distinct roles of entropy modeling and semantic processing—where the former estimates low-dimensional distribution parameters and the latter represents high-dimensional visual contents—we introduce an additional semantic decoder for hyperprior to better address their different requirements. It extracts semantic guidance $c=h_{sem}(\hat{z})$ from the quantized hyperprior, which is injected into the cross-attention layers of the one-step diffusion generator. The diffusion generator refines the initial latent in only one step: $\tilde{y}_{0}=\epsilon_{\theta}(\tilde{y}_{t},c)$, which is then decoded by a pretrained VAE decoder to produce the final reconstruction $\hat{x}=D(\tilde{y}_{0})$.
@@ 3.2.1 From Text to Hyperprior
In the one-step diffusion setting, semantic guidance (the input of the cross-attention layers [44]) plays an even more critical role than in conventional multi-step diffusion. Unlike multi-step methods, which iteratively refine the output and can gradually correct semantic inconsistencies, one-step diffusion models rely entirely on a single forward pass—making accurate semantic conditioning essential. While existing diffusion models typically rely on textual prompts to provide high-level semantic guidance, such prompts are suboptimal in the context of natural image compression. Text struggles to capture localized semantics and often fails to describe fine-grained visual attributes—such as object boundaries and textures—particularly in high-resolution imagery. In addition, generating text descriptions typically requires large models [27], introducing significant computational overhead.
Recent work [18] suggests that the categorical distributed hyperprior with vector quantization (VQ) can capture semantic information. Building on this insight, OneDC replaces textual prompts with categorical hyperprior $\hat{z}$ as semantic guidance for the diffusion model. Specifically, we adopt finite scalar quantization (FSQ) [37], a simpler and more effective quantization scheme than VQ [12], to learn a categorical $\hat{z}$ with semantic information. The hyperprior $z$ is compressed into 7 channels with 4 quantization levels each, yielding an equivalent codebook size of 16,384. At a 64×64$\times$64 × spatial downsampling, this design results in a bitrate of just 0.0034 bpp, incurring very low transmission overhead. Our method offers three key advantages: (1) As shown in Fig. 2, conditioning the one-step diffusion model on categorical hyperprior tokens yields more accurate reconstructions than using text prompts, validating the effectiveness of hyperprior-derived semantic guidance. (2) The 64$\times$ downsampled hyperprior has a large receptive field while preserving spatial locality, effectively balancing global-local semantic expression. (3) Unlike separately trained text encoders [42], our approach enables end-to-end optimization of semantic guidance, allowing the hyperprior to adapt jointly with the diffusion model for improved content alignment.
@@ 3.2.2 Semantic Distillation for Hyperprior
Figure 5: Rate-distortion comparisons. See supplementary for more results.
@@ 3.3 Training Strategy
To enhance training efficiency and stability, we adopt a two-stage training strategy, inspired by prior works [36, 39]. As shown in Fig. 4, the stage I mainly trains the compression module while the stage II focuses on fine-tuning the one-step diffusion model for better reconstruction quality. In both stages, the one-step diffusion model is adapted via LoRA layers [17], allowing fast convergence for the compression task while preserving rich generative priors.
@@
$L_{stageI}=L_{recon}+\lambda R+\alpha L_{aux},\;\;\;\;\text{where}\;\;L_{recon}=L_{1}+L_{perceptual}$ (2)
@@
Stage II: Hybrid-domain Perceptual Learning. This stage fine-tunes the one-step diffusion model to enhance perceptual realism without degrading fidelity. The latent compression module is fixed to improve training stability [36]. Prior work has shown that pixel-level optimization alone is insufficient for optimal perceptual quality [18]. Thus, we adopt a joint training strategy in both pixel and latent domains. Specifically, we integrate the diffusion-based distillation method from [58], which transfers generation knowledge from a pretrained multi-step diffusion teacher. The training also involves both perceptual supervision from the pixel domain and adversarial alignment in the latent domain, which ensures the fidelity by aligning reconstruction with the original image. The total loss is:
$L_{stageII}=L_{distill}+\beta L_{recon}+\gamma L_{adv},\;\;\;\;\text{where:}$: (3)
@@
Here, $\epsilon_{real}$ and $\epsilon_{fake}$ denote the real and fake score network in the teacher model, and $\tilde{y}^{\prime}_{t}$ is the noised latent at timestep $t$. $Disc$ refers to the discriminator. $\beta$ and $\gamma$ balance the pixel and adversarial losses. Additional implementation details are provided in the supplementary material.

### v2

@@ Abstract
[…]
footnotetext: Naifu Xue and Zhaoyang Jia are visiting students at Microsoft Research Asia.
@@ 1 Introduction
The rapid growth in image data has led to increased storage and transmission costs, heightening the need for efficient, high-quality image compression techniques. While VAE-based learned image compression (LIC) [36, 71] has surpassed traditional codecs (e.g., VVC [8]) in rate-distortion (RD) performance, it often produces blurry details at low bitrates [41] since they are typically optimized for objective distortion rather than human perception [7]. To mitigate this, some approaches [41, 44] focus on optimizing visual quality by incorporating perceptual metrics and adversarial losses to enable generative image compression. Despite these advances, such generative models tend to introduce artifacts under extreme compression, degrading the realism of reconstructed images.
@@
In this paper, we introduce OneDC (One-step Diffusion-based generative Codec), a novel framework for ultra-low bitrate image coding. OneDC integrates a latent compression module with a one-step diffusion generator: the former encodes the image into compact latents, while the latter synthesizes high-frequency details conditioned on the latent. Since the pixel-domain training is insufficient for guaranteeing perceptual quality at low bitrates [22], we adopt a hybrid-domain training strategy. Specifically, we combine a pixel-domain perceptual loss to promote fidelity with a latent-domain diffusion distillation objective [66] to enhance realism. This design allows OneDC to effectively leverage the pretrained diffusion model while achieving a balanced trade-off among compression ratio, perceptual quality, and decoding efficiency, as illustrated in Fig. 1.
Moreover, we further explore the role of semantic guidance (i.e., the input of the cross-attention layers [50]) within our framework. In one-step diffusion, such guidance is essential to compensate for the absence of multi-step refinement (see ablation in Section 4.3). While existing diffusion models have utilized text prompts for conditioning, we argue that text is suboptimal for image compression due to two limitations: (1) natural language struggles to capture fine-grained or localized visual semantics, and (2) generating high-quality captions typically requires large-scale vision-language models (e.g., the large BLIP2 [32] used in [9]), introducing substantial computational overhead. This raises a central question: Can we design more effective semantic representations than text for guiding one-step diffusion-based codecs—without incurring excessive computational cost?
Recent studies have shown that the hyperprior in VAE-based codecs can capture high-level semantic information [22, 48]. Compared to textual prompts, hyperpriors provide more precise and spatially aligned semantic cues, making them well-suited for representing localized semantics in high-resolution images (see Fig.2). This insight motivates us to further enhance the representational capability of the hyperprior. Inspired by generative tokenizers [15], where codebooks encode rich semantic content, we propose hyperprior semantic distillation—a training strategy that transfers semantic knowledge from a pretrained tokenizer to the hyperprior through a code prediction module [72]. As illustrated in Fig.2, this approach results in reconstructions that are more semantically accurate and visually coherent.
[…]
@@ 2 Related Work
VAE-based Learned Image Compression. Learned image compression has made rapid progress in RD performance. Ballé et al. [4] introduced an end-to-end framework with analysis/synthesis transforms and a factorized entropy model. The hyperprior model [5] enhanced entropy modeling by encoding side information (hyperprior) to predict distribution parameters, substantially improving RD performance. Subsequent work combined hyperpriors with spatial context [43] and adopts more expressive distribution models such as Gaussian mixtures [11]. Recent studies further refined transform architectures for stronger nonlinear representation [36] and developed advanced entropy models for more compact symbol coding [23], together surpassing the VVC-Intra baseline [8]. Nevertheless, these models are typically optimized with distortion-oriented objectives such as MSE or MS-SSIM [60], which often over-smooth textures and degrade perceptual quality at low bitrates [41], even with SOTA transforms and entropy models [36, 71, 58].
@@
VAE-based Generative Image Compression. To improve perceptual quality at low bitrates, GAN-based approaches such as Agustsson et al. [2] and HiFiC [41] train VAE codecs to synthesize high-frequency details using perceptual losses [24] and adversarial discriminators. MS-ILLM [44] improves realism through a local likelihood model, while EGIC [27] introduces semantic-aware feedback via segmentation-conditioned discriminators. Other methods enhance compression by integrating generative tokenizers: Mao et al. [40] utilize VQGAN [15] for compression, Xue et al. [63] unify token generation and entropy coding, and Jia et al. [22] propose GLC that improves token compression via transform coding. More recently, dual-branch codecs [39, 64] have been introduced to support more compact and flexible latent representations. While these methods highlight the strong feature learning capacity of VAE-based codecs, their reconstruction quality at extremely low bitrates remains limited by model capacity and training scale.
@@
Diffusion-based Generative Image Compression. Diffusion models have rapidly evolved in recent years, progressing from theoretical foundations [51, 20] to practical high-quality implementations [50]. Building on these advances, recent studies [29, 9, 33, 57, 62] have explored the use of generative priors from large-scale pretrained diffusion models in image compression, demonstrating improved perceptual realism compared to conventional VAE-based approaches. Eric et al. [29] combine pretrained diffusion with ControlNet [69] to reconstruct images from captions and compressed sketches. PerCo [9] fine-tunes diffusion models using captions and quantized features, with captions generated by a large BLIP2 model [32]. DiffEIC [33] conditions diffusion models on VAE latents, showing strong performance even without text, and their subsequent work [34] improves encoder and introduces relay fine-tuning. DiffC [57] demonstrates zero-shot compression with diffusion models using reverse-channel coding. DDCM [46] replaces the continuous Gaussian noise in sampling with discrete codebook selections for compression task. While these approaches improve realism at ultra-low bitrates, they often struggle to maintain high fidelity. Moreover, multi-step sampling in DDIM [52] introduces substantial decoding overhead, limiting their practicality. By contrast, our OneDC achieves high fidelity and realism with significantly faster decoding via one-step sampling. While prior work [33] suggests high-level semantics (e.g., text) may be optional, we find such guidance remains crucial in our one-step diffusion for image decoding task, as detailed in Section 4.3.
@@
One-step Diffusion Models. To reduce the computational cost, recent efforts have focused on distilling multi-step models into efficient one-step generators. Song et al. [53] introduce Consistency Models, which achieve few-step generation by enforcing self-consistency along diffusion trajectories. Yin et al. [67] propose Distribution Matching Distillation (DMD), which minimizes an approximate KL divergence between real and generated image distributions modeled by multi-step teacher networks. Their subsequent work, DMD2 [66], further improves generation quality by introducing a diffusion-GAN framework and enhanced training strategies. Based on DMD, Song et al. [54] introduce class-specific student models to improve performance across diverse categories. SwiftBrush series [45, 12] adopt a LoRA adaptation [21] for variational score distillation and further optimize diversity by integrating a clamped CLIP loss. Independently, Kang et al. [25] employ perceptual loss supervision and a multi-scale U-Net discriminator for one-step distillation. These models have also been extended to low-level vision tasks such as super-resolution [14, 61, 59, 26] and image restoration [18], highlighting the potential of one-step diffusion beyond generation. Inspired by these advances, we explore its application in generative image compression, a setting that remains underexplored.
@@ 3 Methodology
We propose OneDC, a one-step diffusion-based generative image compression framework that achieves high-quality reconstruction with low-latency decoding under ultra-low bitrate constraints. Section 3.1 introduces the coding pipeline of our method, consisting of a latent compression module that encodes the input image into a compact latent representation and a one-step diffusion generator that synthesizes reconstructions with realistic details. Given the critical role of semantic guidance in our one-step diffusion and the limitations of textual prompts in compression scenarios, Section 3.2 introduces the use of the hyperprior as an alternative conditioning signal. To further tap the potential of the hyperprior, we propose a semantic distillation strategy that transfers knowledge from a pretrained generative tokenizer to the hyperprior codec, thereby improving semantic accuracy. Finally, Section 3.3 presents our training scheme, which combines pixel-domain and latent-domain objectives to jointly optimize reconstruction fidelity and perceptual realism.
@@ 3.1 Framework Overview
[…]
@@
[…]
@@ 3.2.1 From Text to Hyperprior
In the one-step diffusion setting, semantic guidance (the input of the cross-attention layers [50]) plays an even more critical role than in conventional multi-step diffusion. Unlike multi-step methods, which iteratively refine the output and can gradually correct semantic inconsistencies, one-step diffusion models rely entirely on a single forward pass, making accurate semantic conditioning essential. While existing diffusion models typically rely on textual prompts to provide high-level semantic guidance, such prompts are suboptimal in the context of natural image compression. Text struggles to capture localized semantics and often fails to describe fine-grained visual attributes—such as object boundaries and textures—particularly in high-resolution imagery. In addition, generating text descriptions typically requires large models [32], introducing significant computational overhead.
Recent work [22] suggests that the categorical distributed hyperprior with vector quantization (VQ) can capture semantic information. Building on this insight, OneDC replaces textual prompts with the categorical hyperprior $\hat{z}$ as semantic guidance for the diffusion model. To compress the hyperprior, we adopt finite scalar quantization (FSQ) [42], a simpler yet more effective alternative to VQ [15], to learn a categorical $\hat{z}$ with semantic information. The hyperprior $z$ is compressed into 7 channels with 4 quantization levels each, yielding an equivalent codebook size of 16,384. At a $64\times$ spatial downsampling, this design results in a bitrate of just 0.0034 bpp, incurring very low transmission overhead. To substitute for text embeddings, we introduce a semantic decoder $h_{sem}$ that transforms the quantized hyperprior $\hat{z}$ into a semantic context $c\in\mathbb{R}^{B\times N\times D}$, where $N=H^{\prime}\times W^{\prime}$ is the flattened spatial shape and $D$ is the embedding dimension. This context is injected into the cross-attention layers of the one-step diffusion U-Net, replacing the textual embeddings used in the original design. In each cross-attention layer, the input latent feature $f_{in}$ serves as the query, while the semantic context $c$ provides both keys and values:
$f_{out}=\mathrm{Softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)V,\quad\text{where}~Q=W_{Q}f_{in},~K=W_{k}c,~V=W_{v}c$ (1)
where $d_{k}$ denotes the dimensionality of the key vectors and $f_{out}$ is the output feature of the cross-attention layer. This design enables every spatial location in the latent to attend adaptively to semantically relevant tokens, strengthening content alignment. Compared with text conditioning, hyperprior-based guidance offers three advantages: (1) As shown in Fig. 2, conditioning the one-step diffusion model on categorical hyperprior tokens yields more faithful reconstructions than using text prompts, validating their effectiveness as semantic guidance. (2) The $64\times$ downsampled hyperprior provides a large receptive field while retaining spatial locality, thereby offering a more balanced global–local semantics compared to purely global text embeddings. (3) Unlike separately trained text encoders [49], our approach enables end-to-end optimization of semantic guidance, allowing the hyperprior to adapt jointly with the diffusion model for improved content expression.
@@ 3.3 Training Strategy
To enhance training efficiency and stability, we adopt a two-stage training strategy, inspired by prior works [41, 44]. As shown in Fig. 4, stage I mainly trains the compression module while stage II focuses on fine-tuning the one-step diffusion model for better reconstruction quality. In both stages, the one-step diffusion model is adapted via LoRA layers [21], allowing fast convergence for the compression task while preserving rich generative priors.
@@
[…]
@@
Figure 7: Rate–distortion curves under the full-resolution setting. Zoom in for better view.
Figure 8: Rate–distortion curves under the resize & center-crop on Kodak $\times512$ and CLIC2020 $\times768$. Zoom in for better view.
Stage II: Hybrid-domain Perceptual Learning. This stage fine-tunes the one-step diffusion model to enhance perceptual realism without degrading fidelity. The latent compression module is fixed to improve training stability [41]. Prior work has shown that pixel-level optimization alone is insufficient for optimal perceptual quality [22]. Thus, we adopt a joint training strategy in both pixel and latent domains. Specifically, we integrate the diffusion-based distillation method from [66], which transfers generation knowledge from a pretrained multi-step diffusion teacher. The training also involves both perceptual supervision from the pixel domain and adversarial alignment in the latent domain, which ensures fidelity by aligning reconstruction with the original image. The total loss is:
$L_{stageII}=L_{distill}+\beta L_{recon}+\gamma L_{adv},\;\;\;\;\text{where:}$ (4)
@@
Here, $\epsilon_{real}$ and $\epsilon_{fake}$ denote the real and fake score networks in the teacher model, and $\tilde{y}^{\prime}_{t}$ is the noised latent at timestep $t$. DiscDisc refers to the discriminator. $\beta$ and $\gamma$ balance the pixel and adversarial losses. Additional implementation and loss details are provided in the supplementary material.